OpenStudy (deleonbrrj):

how do I fix this? 2/5 * (5)^5/2

OpenStudy (amistre64):

how did you break it?

OpenStudy (ash2326):

\[\large \frac25 \times 5^{\frac52}\] Bring the 5 from the denominator to the numerator \[\large \frac{2}{\underline {5}} \times 5^{\frac52}\] \[\large2 \times 5^{\frac52}\times 5^{-1}\] \[\large2 \times 5^{\frac52-1}\] \[\large2 \times 5^{\frac52-\frac22}\] Finally we get \[\Huge2 \times 5^{\frac32}\]

OpenStudy (deleonbrrj):

no, is integration \[\int\limits_{0}^{5}x ^{3/2}\]

OpenStudy (amistre64):

power rule for integration: \[\int x^n dx=\frac{1}{n+1}x^{n+1}\]

OpenStudy (deleonbrrj):

perfect ash2326

OpenStudy (deleonbrrj):

i I knew the result but did not know how to get it, thanks!

OpenStudy (ash2326):

How did I help? I misinterpreted the question!!!

OpenStudy (amistre64):

you simplified the results :)